OpenStudy (anonymous):

(x+3^(1/3()(x-3^(1/3))/(x-27) as limit x approaches 27.find the limit.

OpenStudy (turingtest):

\[\lim_{x\to27}\frac {(x+3^{1/3})(x-3^{1/3})}{x-27}\]correct?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

are you sure you typed that correctly? none of the x's are squared or cubed or anything?

OpenStudy (anonymous):

no.i misprint it .The 1/3 will be over x,not over 3.

OpenStudy (turingtest):

that makes much more sense\[\lim_{x\to27}\frac {(x^{1/3}+3)(x^{1/3}-3)}{x-27}\]

OpenStudy (turingtest):

factor the denominator using difference of cubes

OpenStudy (turingtest):

\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

yes i am doing.

OpenStudy (turingtest):

ok, let me know if/where you get stuck

OpenStudy (anonymous):

yes,i got the answer 2/9,thank you very much.

OpenStudy (turingtest):

you're welcome :)